# ELTEX: A Framework for Domain-Driven Synthetic Data Generation

### v1

@@ 4.2 Task Complexity
4.3 Fine-Tuning and Deployment
We fine-tuned Gemma-2B using PEFT Mangrulkar et al. (2022) via LoRA Hu et al. (2022) while maintaining high computational efficiency through quantization (see Appendix C for full fine-tuning and deployment specifications).
5 Results
This section presents our evaluation of various models on the cybersecurity implication detection task, where each model classifies social media messages for potential blockchain attack relevance using a 0-1 scoring system. We first analyze overall performance across standard metrics, then examine the effectiveness of synthetic data, and finally provide a comparative analysis with models designed for different objectives.
Table 3 summarizes the performance metrics for all evaluated models: the base Gemma-2B model, fine-tuned variants (on real data, synthetic data, and a hybrid approach), and comparative benchmarks including GPT-4o, the general-purpose reasoning model Granite-3.2-2B IBM (2025), and cybersecurity-focused LLMs fine-tuned on the Primus dataset Yu et al. (2025).
Model Performance Metrics Error Rates
Acc. Brier Recall F1 ROC False Pos. False Neg.
Gemma-2b-it 0.51 0.43 0.30 0.30 0.44 0.24 0.25
Gemma-2b-Real 0.65 0.31 0.47 0.61 0.73 0.03 0.32
Gemma-2b-Synth 0.77 0.16 0.78 0.76 0.85 0.13 0.11
Gemma-2b-Hybrid 0.82 0.14 0.79 0.81 0.88 0.08 0.10
Granite-3.2-2b-Instruct 0.74 0.39 0.58 0.68 0.73 0.06 0.20
Llama-Primus-Instruct 0.71 0.29 0.43 0.58 0.69 0.03 0.27
Llama-Primus-Merged 0.76 0.23 0.59 0.70 0.76 0.04 0.19
GPT-4o 0.84 0.10 0.71 0.81 0.94 0.02 0.14
Table 3: Performance on social media cyberattack detection. Metrics reflect model ability to score messages (0-1) for blockchain attack relevance. Best scores in bold, second-best underlined.
5.1 Overall Performance Analysis
Our results show significant performance improvements through fine-tuning, with the hybrid approach (combining synthetic and real data) achieving the best results among our Gemma-2B variants. The base Gemma-2B model shows limited effectiveness (0.51 accuracy, 0.30 F1), while the hybrid model achieves performance (0.82 accuracy, 0.81 F1) competitive with the much larger GPT-4o (0.84 accuracy, 0.81 F1).
The Brier score, which measures both calibration and accuracy of probabilistic predictions (lower is better), shows that synthetic-trained models (0.14-0.16) are better calibrated than other variants in our evaluation (0.23-0.43). This suggests that our synthetic data approach improves classification accuracy and helps models produce more reliable confidence estimates, which is crucial for applications requiring well-calibrated risk assessments. For context, the Brier score represents the mean squared difference between predicted probabilities and actual outcomes, with perfect predictions scoring 0 and worst possible predictions scoring 1.
5.2 Synthetic Data Effectiveness
Domain Adaptation
The substantial improvement from the base Gemma-2b (0.51 accuracy) to the hybrid model (0.82 accuracy) demonstrates synthetic data’s value for adapting models to informal cybersecurity discourse. Our ELTEX-generated examples appear to help models recognize patterns characteristic of social media discussions about blockchain security, including veiled threats and technical jargon that may indicate potential attack vectors.
Error Profile Analysis
The different error profiles across training approaches reveal important trade-offs. Models trained solely on real data show low false positive rates (0.03) but miss many actual threats (0.32 false negative rate). In contrast, synthetic-trained models show improved threat detection (0.10-0.11 false negative rates) at the expense of slightly higher false positives (0.08-0.13). This suggests synthetic data helps capture a wider range of threat expressions, potentially beneficial for early warning systems where failing to detect threats carries higher costs than false alarms.
Complementary Training Effects
The hybrid approach combining synthetic and real data achieves the best overall performance (0.82 accuracy, 0.14 Brier score), outperforming both individual training approaches. This finding aligns with recent research by Li et al. (2024), who observed that combining synthetic and real data often improves both performance and training stability. The complementary nature of these data sources likely helps the model generalize across both common and rare threat expressions while maintaining reasonable precision.
5.3 Comparative Analysis with Different Model Types
We include results from models with different design objectives to contextualize our approach within the broader language model landscape. These comparisons are not intended as direct competitive benchmarks but rather to understand how different model architectures and training approaches perform on this specialized task.
GPT-4o serves as a performance ceiling (0.84 accuracy, 0.10 Brier score), demonstrating what’s achievable with a state-of-the-art large-scale model. Notably, our hybrid Gemma-2B model achieves the same F1 score (0.81) as GPT-4o despite having significantly fewer parameters, suggesting that targeted fine-tuning can partially compensate for model scale on specialized tasks.
The cybersecurity-focused Primus models (0.71-0.76 accuracy) provide context on how domain-specific training transfers to our particular task. While these models show strong general cybersecurity knowledge, their performance on social media threat detection suggests that domain knowledge alone is insufficient without task-specific tuning.
Similarly, Granite-3.2-2B (0.74 accuracy, 0.39 Brier score), known for cybersecurity capabilities, performs reasonably well on classification accuracy but shows poorer calibration. This indicates that general cybersecurity abilities provide some transfer to specialized tasks, but well-calibrated confidence estimates may require more targeted training.
These results suggest that for specialized applications like blockchain threat detection in social media, a combination of targeted synthetic data generation and fine-tuning can enable smaller models to achieve performance comparable to much larger general-purpose models on key metrics.
@@ 6 Conclusion
We introduced ELTEX, a domain-driven framework for generating synthetic training data that enables effective adaptation of smaller language models to specialized domains.
Our results highlight three key findings: (1) ELTEX’s systematic token extraction approach produces high-quality synthetic data that captures diverse attack patterns while maintaining domain validity, (2) the integration of domain-specific indicators in synthetic data generation leads to better coverage of potential threats, reducing false negatives without significantly increasing false positives, and (3) our domain-driven approach enables resource-efficient models to achieve performance competitive with larger architectures while maintaining strong calibration.
Limitations
While ELTEX demonstrates promising results in domain adaptation for cybersecurity and offers unique advantages through synthetic data generation — particularly in providing comprehensive topic coverage and modernizing historical data — several important limitations warrant discussion:
Data Dependencies and Scalability
ELTEX requires an initial corpus of high-quality, real-world data to seed the synthetic data generation process. Consequently, the quality and diversity of the synthetic output are inherently tied to the representativeness of this seed data, which may limit ELTEX’s applicability in domains where even minimal real data is challenging to obtain or validate. Moreover, while our experiments confirm ELTEX’s effectiveness at moderate scales, its performance and computational efficiency have not been validated on significantly larger datasets (e.g., exceeding 100,000 samples). The multi-stage token extraction and deduplication processes could exhibit non-linear increases in computational cost, potentially hindering scalability in high-volume applications.
Model Dependencies and Reproducibility
The current implementation of ELTEX heavily leverages GPT-4o for synthetic data generation. This introduces several concerns:
  • Version Dependencies: Updates or changes in GPT-4o could alter generation patterns, complicating reproducibility and longitudinal studies.
  • Access Constraints and Costs: API access limitations and associated costs may restrict the framework’s broader adoption, particularly in resource-constrained settings.
  • Propagation of Biases: Any inherent biases or limitations present in GPT-4o can be transferred to the synthetic data, potentially affecting downstream model performance.
  • Classification Score Granularity: While GPT-4o generates continuous classification scores (0 to 1), our human verification process only validates binary labels. This discrepancy means we may lose potentially valuable nuance in the model’s confidence levels, and the relationship between continuous scores and final binary classifications warrants further investigation.
Although ELTEX is designed to be model-agnostic, our current evaluation has not systematically compared alternative large language models, leaving open questions regarding the generalizability of the synthetic data quality across different generation backends.
QA Scalability
The human-in-the-loop QA step plays a critical role in maintaining data integrity. However, this process poses significant scalability challenges as the dataset size increases. Even with the assistance of tools like NotebookLM, manual verification remains resource-intensive and demands substantial domain expertise. This bottleneck may impede rapid dataset expansion and could lead to inconsistent QA outcomes when applied across different subdomains or technical areas.
Evaluation
Our evaluation focuses on a specific aspect of cybersecurity, detecting potential blockchain attack signals in social media messages. This task represents just one facet of the broader cybersecurity domain, and model performance should be interpreted within this limited context. The Primus models, being developed for general cybersecurity applications, may offer advantages in areas not measured by our specific task. Their different performance characteristics on our social media task should not be interpreted as reflecting their overall cybersecurity capabilities. Similarly, Granite’s performance without task-specific tuning may indicate how general-purpose reasoning models can be applied to specialized security monitoring tasks, rather than representing their full capabilities. These results suggest that synthetic data generation may be a useful approach for adapting models to specific cybersecurity detection tasks, particularly in domains where labeled data is limited.
Domain Generalization
While our case study validates ELTEX in the context of cybersecurity in blockchain, its generalizability to other specialized domains remains untested. Adapting the token extraction and dynamic prompt construction methodologies to domains with different linguistic characteristics or technical complexities may require substantial modifications. Future work should explore these adaptations and evaluate the framework’s performance across a broader range of domains.
These limitations highlight several promising avenues for future research, including the exploration of alternative or hybrid data sources, the development of more scalable and automated QA methodologies, and rigorous testing of ELTEX’s applicability and performance in diverse specialized domains.
@@
All data collection procedures were conducted in accordance with X’s official API (v2) terms of service and privacy policies X (2025). We ensured that our data acquisition methods did not violate any platform-specific rules or regulations.
@@ B.2 Human Annotation Procedure
Human annotators used the preliminary labels generated by GPT-4o (version August 6, 2024) as a starting point for the annotation process. Guided by the list of cyberattack indicators, they reviewed the assigned cyberattack_score and adjusted the final labels accordingly. This approach leverages the efficiency of LLM-generated insights while ensuring high annotation quality.
@@ B.3 Validation and Accuracy
Validation confirmed that GPT-4o (version August 6, 2024) achieved over 95% accuracy.
@@ D.2.1 Retention Trends
D.3 Clustering Analysis on Synthetic Data
We performed semantic clustering analysis on the BGE-base-en-v1.5 embeddings Xiao et al. (2023) of both real and synthetic messages. The technical details of our implementation include:
  • Clustering Algorithm: k-means MacQueen (1967) with k = 10 and k-means++ initialization (random_state = 42)
  • Dimensionality Reduction: UMAP visualization with parameters:
    • n_components = 2
    • metric = cosine
    • random_state = 42
  • Dataset Sizes:
    • Real cyberattack messages: 994
    • Synthetic cyberattack messages: 6,941
(a) Real Data
(b) Synthetic Data
Figure 4: UMAP Visualization of K-means Clustering Results
The UMAP visualization (Figure 4) reveals different clustering patterns between real and synthetic data. Real data shows some well-separated clusters alongside overlapping ones, while synthetic data displays more uniform overlap and elongated structures. This suggests that real data contains both distinct security events and related incident types, whereas synthetic data forms a more continuous semantic space.
Tables 5 and 6 present the most representative clusters from each dataset, highlighting their distinct semantic organization.
Cluster Top Keywords (frequency)
Euler Finance finance (129), million (89), defi (60), protocol (59), exploit (55), attack (38), euler (37), exploited (35), curve (32)
Poloniex/HTX poloniex (63), exchange (58), htx (53), million (42), justin (32), sun (32), crypto (30), hack (30), hacker (24)
Bridge Exploit bridge (51), multichain (40), orbit (38), chain (31), million (29), hack (27), exploit (18), protocol (14), hacked (11)
Atomic Wallet wallet (73), atomic (68), users (21), hack (19), security (14), funds (11), crypto (10), million (9)
Table 5: Representative Clusters in Real Data (Event-Specific Organization)
Cluster Top Keywords (frequency)
Network transaction (366), attack (274), network (269), potential (260), mining (250), blockchain (216), unusual (202), alert (174)
Wallet dormant (245), wallet (96), addresses (81), activation (77), activity (73), suddenly (62), sudden (54), wallets (53)
Market price (368), manipulation (230), market (207), exchange (172), rapid (171), fluctuations (90), trading (85)
Governance governance (333), community (206), sentiment (86), irregularities (79), proposal (73), manipulation (63), voting (49)
Table 6: Representative Clusters in Synthetic Data (Category-Based Organization)
The keyword analysis shows that real data clusters form around specific security incidents (e.g., the Euler Finance exploit), characterized by concrete terms ("million", "hack") and named entities. In contrast, synthetic data clusters represent broader cyberattack categories (e.g., network anomalies, market manipulation), employing more general terminology ("potential", "unusual", "alert").
This pattern aligns with our generation approach: synthetic messages were created to cover a broad spectrum of potential indicators rather than specific incidents. While this results in wider coverage of possible cyberattack scenarios, it comes at the cost of reduced event-specific detail. This trade-off suggests that combining real and synthetic data might provide both concrete incident coverage and broad indicator diversity.
@@ Appendix F Costs Analysis
Our synthetic data generation relied on the GPT-4o (version August 6, 2024) model served via Azure OpenAI. The pricing structure was as follows:

### v2

@@ 5.1 Task Complexity
5.2 Dataset Construction
Our training pipeline leverages the dataset of 1,603 real-world social media messages, which we first split into training (80%) and validation (20%) sets. Using ELTEX, we expanded only the training portion into a synthetic dataset of 8,892 messages (5,530 cyberattack, 3,362 general), with generated classification scores ranging from 0 to 1. We collected an independent test set of 398 real-world messages (206 labeled as cyberattack and 192 as general), drawn from entirely distinct attack events not present in our training data. Table 3 in Appendix C shows the distribution of primary attack types across training and test sets.
5.3 Key Findings
Our evaluation revealed three critical insights for industrial deployment:
Enhanced performance with resource efficiency:
Fine-tuning Gemma-2B with ELTEX-generated data achieved classification performance (F1: 0.81) comparable to GPT-4o while reducing inference costs. This represents significant operational savings for high-volume monitoring scenarios typical in security operations.
Model calibration:
ELTEX-enhanced models demonstrated substantially better calibration (Brier score: 0.14-0.16) than both the base model (0.43) and domain-specialized LLMs, like Primus (0.23-0.29). Well-calibrated confidence scores enable more effective alert prioritization and risk assessment in production environments.
Balanced error profiles:
Our hybrid approach (real data + ELTEX) achieved a favorable balance between false positives (0.08) and false negatives (0.10), unlike models trained solely on real data which showed a strong bias toward one error type (0.03 and 0.32 respectively). This balanced performance enables more reliable threat detection for operational security contexts where both error types carry costs. Notably, the hybrid approach achieved the best overall performance, suggesting complementary benefits from both data sources, which aligns with recent research by Li et al. (2024).
@@ 6 Conclusion
ELTEX addresses domain specialization challenges by systematically preserving domain knowledge throughout synthetic data generation. Our cybersecurity case study demonstrated three key advantages: comparable performance to frontier LLMs with significantly reduced computational requirements, superior model calibration, and balanced error profiles critical for operational contexts. The Google Sheets implementation democratizes domain adaptation for non-technical users. Limitations include reliance on seed data for generation and validation limited to a classification task in one domain.
@@
All data collection procedures were conducted in accordance with X’s official API (v2) terms of service and privacy policies X (2025). We ensured that our data acquisition methods did not violate platform-specific rules or regulations.
@@
Google Sheets Implementation
Our Google Sheets add-on stores user data solely within the user’s spreadsheet, ensuring direct control over content. The add-on minimally processes user email addresses — provided by Apps Script login — for operational purposes (e.g., linking requests to the correct spreadsheet). In line with best practices for personally identifiable information (PII), we neither persist these emails beyond short-term workflows nor share them externally. Users remain responsible for ensuring that any data they enter complies with applicable privacy and ethical guidelines. Meanwhile, embeddings used by the deduplication step are retained externally for up to 24 hours, and then removed. This design respects data ownership and privacy while enabling efficient collaboration and real-time data generation.
@@ B.2 Human Annotation Procedure
Human annotators used the preliminary labels generated by GPT-4o as a starting point for the annotation process. Guided by the list of cyberattack indicators, they reviewed the assigned cyberattack_score and adjusted the final labels accordingly. This approach leverages the efficiency of LLM-generated insights while ensuring high annotation quality.
@@ B.3 Validation and Accuracy
Validation confirmed that GPT-4o achieved over 95% accuracy.
@@ Appendix F Costs Analysis
Our synthetic data generation relied on the GPT-4o model served via Azure OpenAI. The pricing structure was as follows: